### v1

# When Algorithms Meet Artists: Topic Modeling the AI-Art Debate, 2013–2025

Abstract
As generative AI continues to reshape artistic production and alternate modes of human expression, artists whose livelihoods are most directly affected have raised urgent concerns about consent, transparency, and the future of creative labor [9, 1]. However, the voices of artists are often marginalized in dominant public and scholarly discourse [2]. This study presents a twelve-year analysis, from 2013 to 2025, of English-language discourse surrounding AI-generated art. It draws from 439 curated 500-word excerpts sampled from opinion articles, news reports, blogs, legal filings, and spoken-word transcripts. Through a reproducible methodology, we identify five stable thematic clusters and uncover a misalignment between artists’ perceptions and prevailing media narratives. Our findings highlight how the use of technical jargon can function as a subtle form of gatekeeping, often sidelining the very issues artists deem most urgent. Our work provides a BERTopic-based methodology [6] and a multimodal baseline for future research, alongside a clear call for deeper, transparency-driven engagement with artist perspectives in the evolving AI-creative landscape.
Keywords: Generative AI • Art • Topic Modeling • BERTopic • UMAP • Transformer Embeddings
1 Introduction
Since the introduction of generative adversarial networks (GANs) in 2014  [4] and the development of the transformer architecture in 2017 [19], generative AI has accelerated from a niche deep learning community to a mass market tool. Milestones have since followed in quick succession: Google’s DeepDream popularised neural‑style image generation [11]; OpenAI’s DALL-E demonstrated large-scaled text-to-image synthesis [13]; and in 2022 the release of Stable Diffusion [15] and Midjourney dramatically lowered the barrier to entry, unleashing a flood of AI-generated imagery across platforms. Thus far, each technical breakthrough has intensified public debate over authorship, labor, consent, attribution, and fair compensation. These debates are now playing out in the court through class-action lawsuits (e.g., Andersen v. Stability AI, 2023 [1]) and in emerging regulation such as the EU AI Act [3].
Artists have not been passive observers of this shift. Survey research by Lovato et al. [9] reveals that many practitioners regard non‑consensual dataset scraping as a serious threat and advocate for full transparency around AI training inputs. This survey identifies four dominant frames in practitioners’ own words:
RQ1 Threat:
Do artists perceive generative AI as a material threat to their livelihoods?
RQ2 Utility:
Do they also recognise practical or creative benefits?
RQ3 Transparency:
Do they demand detailed disclosure of training data and model provenance?
RQ4 Ownership:
How do they assign rights among original artists, model developers, and end users?
While Lovato et al. provides a space for artists to voice concern over non-consensual scraping and call for full dataset transparency, it remains unclear to what extent public discourse foregrounds or overshadows these artist-identified concerns. Prior computational studies have primarily focused on short‑form content such as social media [8] or formal scholarly literature [17], leaving the broader, multimodal public discourse on generative AI and art largely unmapped. Our study builds on this literature by offering the first multimodal, consensus-based map of generative AI art discourse, identifying both clear topic categories and overlapping themes that appropriately interpret public conversation. Specifically, in this paper, we analyze twelve years (2013–2025) of English-language discourse across blogs, podcasts, legal documents, news articles, and academic papers. Through leveraging transformer‑based embeddings, k‑means clustering, and BERTopic [6], together with consensus‑based dimensional reduction, we aim to address the following central questions:
  • Topic structure: What coherent clusters structure public discourse on generative AI art?
  • Artist concerns: How do these clusters relate to artist-identified themes Threat, Utility, Transparency, and Ownership [9]?
  • Temporal dynamics: How have these themes shifted in prominence over time, particularly around major technical releases or legal events?
We hypothesize that public discourse partially aligns with the four artists’ frames (Threat, Utility, Transparency, Ownership) identified by Lovato et al. [9]. Specifically, technical narratives will dominate and drown out artists’ concerns of Transparency during major model releases.
Contributions
We identify five thematically distinct topics and situate them along two continuous axes (D1: Artistic to Technical Engineering; D2: General Market to Legal & Regulatory).
We quantify year-by-year shifts in cluster prevalence against major generative-AI milestones (2014–2025).
We map those empirical clusters back to artist perceptions, revealing where public discourse amplifies versus overshadows artist voices.
2 Methods
2.1 Data Collection and Preprocessing
We assembled a multimodal corpus of English-language discourse on AI-generated art spanning from 2013 through early 2025. Using a curated set of search terms (appendix) we gathered content from a range of sources, including academic journals, conference proceedings, policy briefs, legal rulings, blog posts, opinion pieces, podcasts, YouTube, TEDx presentations, and artist interviews. All spoken media were transcribed using automated speech recognition (ASR) and manually reviewed for accuracy. Each source was tagged by media type (e.g., article, podcast, peer-reviewed paper) and categorized by subgenre (e.g., legal commentary, artist panel, solo talk) to support metadata-aware analysis.
To prepare the data for natural language processing, we applied a standardized preprocessing pipeline. This included lowercasing, removing URLs, dates, numerals, and formatting artifacts, followed by lemmatization and stop-word removal. To minimize selection bias, we also excluded search query terms such as “AI,” “art,” “artwork,” and “technology.” Each document was segmented into non-overlapping 500-word sections, producing a final dataset of 439 text chunks of relatively uniform length for downstream embedding and analysis.
2.2 Embedding, Clustering, and Topic Modeling
We generated 768-dimensional sentence embeddings for each chunk using the all-mpnet-base-v2 model from SentenceTransformers [14]. To ensure stable clustering, we applied L2-normalization to all embeddings.
We then used a hybrid topic modeling pipeline that combines k-means clustering with BERTopic [6], which augments centroid assignments with class-based TF-IDF representations [18]. This approach leverages the semantic nuance of transformer embeddings while yielding coherent and readable topic descriptions.
2.3 Cluster Selection and Stability
Recognizing that k-means clustering can vary based on initialization, we ran the model across five random seeds (15, 158, 24, 5, and 336), each selected from a uniform distribution between 1 and 1000. For each random seed we tested k-means solutions with K=2K=22020, recording the within-cluster sum of squares (inertia), Silhouette scores [16], and the pairwise Adjusted Rand Index (ARI; [7]). Although the Silhouette metric reached its maximum at K=2K=2–3 (max $\approx 0.39$ for one seed), these very small $K$ values collapsed several distinct narratives into single clusters. We therefore adopted $K=5$: it obtained the second-highest Silhouette score for three of the four seeds, produced strong inter-seed agreement (mean ARI = 0.63, versus 0.71 for $K=4$), and offered the most interpretable clusters spanning technical, legal, community, creative, and economic themes.
2.4 Consensus Dimensions via Procrustes-Aligned UMAP
To extract continuous dimensions of discourse, we projected each BERTopic output into two dimensions using UMAP [10] with fixed random states. We then aligned the five UMAP outputs to a common reference using Procrustes analysis [5]. This alignment process showed minimal variation across embeddings, with disparities ranging from 0.025 to 0.045, indicating strong geometric consistency. By averaging across aligned outputs, we derived two consensus axes: $D1_{consensus}$ and $D2_{consensus}$. These capture key semantic trends. The first axis ranges from community-oriented, artistic discourse to technical, engineering language. The second axis spans from conversational, general-market framing to formal legal and intellectual property concerns.
2.5 Downstream Analyses
With both discrete topic labels and continuous axis coordinates in place, we conducted several downstream analyses. First, we tracked how the prevalence of each topic evolved over time, focusing on the period from 2013 to 2025. This temporal mapping was annotated with major generative AI milestones such as the launch of DALL·E, Stable Diffusion, and significant legal developments including Andersen v. Stability AI.
We augmented every section with additional linguistic features, including sentiment, lexical diversity, and word count, using DistilBERT-SST2 [20], the twitter-roberta-sentiment model, TextBlob polarity and subjectivity scores, and readability indices from the textstat library. These metrics were then correlated with both topic membership and positions on the $D1$ and $D2$ semantic axes. To ground our qualitative interpretations, we extracted the twenty most frequent unigrams and bigrams in the uppermost and lowermost five percentiles of $D1$ and $D2$, thereby isolating vocabulary unique to each rhetorical extreme. The resulting section-level correlations are visualised in Figure 5 and summarised in Table 3.
2.6 Statistical Alignment with Artist Survey
To quantify how closely public discourse mirrors the four artist-identified frames from Lovato et al. (2024) [Threat (RQ1), Utility (RQ2), Transparency (RQ3), and Ownership (RQ4)], we transformed each section’s consensus coordinates ($D1$, $D2$) into four nonnegative “alignment weights.”
We began by linearly rescaling each axis to [1,1][-1,1]. From there, we then computed raw quadrant scores by multiplying the appropriate half-axis projections (e.g. Threat == (D1)times(D2)(-D1)\\ times(-D2)), and normalized those so that for each section the four weights sum to one. This yields per-section alignment scores: $p_{threat}$, $p_{utility}$, $p_{transparency}$, $p_{ownership}$.
At the corpus level, we compared the mean observed alignment proportions
$\hat{P}=(\bar{p}_{threat},\bar{p}_{utility},\bar{p}_{transparency},\bar{p}_{ownership})$ (1)
against the Lovato survey benchmarks (after normalizing the original values):
$P_{0}^{original}=(0.619,0.449,0.802,0.414)$ (2)
$P_{0}=(0.271,0.197,0.351,0.181)$ (3)
We then performed one-sample z-tests for proportions using the proportions_ztest function from the Statsmodels package to determine whether observed alignment scores differed significantly from the survey-derived expectations.
3 Results
3.1 Consensus Axes
Figure 1: Two‐dimensional consensus UMAP embedding with $k=5$ centroids and Lovato et al. RQs. The figure is annotated our qualitative interpretation of the D1 and D2 axes by looking at uni- and bi-grams that were exclusive to the top and bottom 5% of data on each axis.
Topic $D1_{centroid}$ $D2_{centroid}$ Quadrant Dominate RQ
0 0.0270 0.0075 I (+D1/+D2) NA
1 0.0119 -0.0479 IV (+D1/-D2) Utility (RQ2)
2 -0.0261 0.0252 II (-D1/+D2) Ownership (RQ4)
3 -0.0216 -0.0246 III (-D1/-D2) Threat (RQ1)
4 -0.0020 0.0627 II (-D1/+D2) Transparency (RQ3)
Table 1: Consensus-topic centroids on $D1/D2$, qualitative region, and Lovato et al. research-question mappings. Quadrant refers to the sign of ($D1,D2$): I = (+, +), II = (-, +), III = (-, -), IV = (+, -). Dominate Lovato et al. Research Questions (RQ) for each topic. Topic 0 is specifically about technical engineering terms (i.e. algorithm details) and was the only cluster of text not represented by Lovato’s et al RQs.
Figure 1 presents the two-dimensional UMAP projection of discourse sections, with cluster centroids and labels overlaid. Each section is colored by its assigned topic, and the axes are annotated with interpretive labels grounded in both qualitative reading and alignment with Lovato et al.’s artist-centered research questions.
The horizontal axis ($D1$) spans from community-based, creative practice language on the left to technical engineering discourse on the right. For example, Topic 2 (“Community & Artistic Practice”) and Topic 3 (“Threat”) are located on the artistic side, while Topic 0 (“Technical Engineering”) and Topic 1 (“Utility & Market”) are positioned on the technical end.
The vertical axis ($D2$) distinguishes between general-market and conversational framing (lower values) and legal or regulatory discourse (higher values). Topic 4 (“Legal & Regulatory”) anchors the upper region of $D2$, reflecting more formal and policy-oriented language. Meanwhile, Topic 1 is situated in the lower-right quadrant, reflecting its mix of optimistic market framing and technical adoption. The position of each topic within this semantic space offers an interpretable map of how public discourse is structured and how artist concerns are distributed across different domains.
3.2 Topics over Time
Figure 2: Proportion of each topic by year (2013–2025), with milestone annotations (light gray).
Year Event
2014 GANs introduced (Goodfellow et al.)
2015 Google DeepDream release
2017 Transformer architecture published (Vaswani et al.)
2018 (*1) AI art auction at Christie’s; BERT release
2019 GPT‑2 release
2020 (*2) GPT‑3 public API; authorship debates
2022 (*3) Stable Diffusion open‑source; GPT‑3.5
2023 (*4) GPT‑4 release; Andersen v. Stability AI lawsuit
2024 GPT‑4o multimodal launch
2025 (*5) EU AI Act negotiations; Christie’s first AI auction
Table 2: Key generative AI and art milestones (2014 - 2025)
Figure 2 shows the annual proportion of topics across text sections from 2013 through 2025, with key generative AI and art milestones overlaid in light gray (Table 2).
Early coverage (2013–15) was dominated by Topic 2 (55% $\rightarrow$ 15%) and Topic 0 (10% $\rightarrow$ 58%), reflecting initial community debates and excitement around DeepDream. With the 2017 Transformer breakthrough and 2018 auction headlines, Topic 2 and Topic 0 surge again ($\approx$50%), while transparency-focused Topic 4 peaks (15% $\rightarrow$ 30%). Post-2020, Topic 1 “Utility/Market” explodes to 45% in 2021 around GPT-3.5 and Stable Diffusion, then stabilizes at roughly one-third share. Topic 3 “Threat” and Topic 2 “Ownership” re-emerge modestly through 2025 (each $\approx$15–25%), but never regain early prominence.
In 2013, Topic 2 (“Community & Artistic Practice”) dominated the discourse—accounting for over 50% of all 500-word sections—followed by Topic 3 (“Threat Narratives”) at just under 30%, while Topics 0 (the furthest cluster/topic on the “Technical Engineering Discourse” end of the $D1$ spectrum) and 1 (the furthest cluster/topic on the “General Market Discourse” end of the $D1$ spectrum) each comprised roughly 10%.
Following the 2015 release of Google DeepDream, we observe a sharp decline in Topic 2 (from 54% to 15% by 2015) concurrent with a surge in Topic 0, which rises to nearly 60% in 2015. Over the next two years, “Technical Engineering Discourse” (Topic 0) recedes to about 25% by 2017, while Topic 4 (the furthest cluster/topic towards the “Legal & Regulatory Discourse” end of the $D2$ spectrum) first appears at 15% in 2015 and peaks at 30% in 2017 before vanishing entirely by 2018.
The 2018 transformer breakthrough and subsequent API launches for GPT-2 and GPT-3 mark another transition. Topic 0 rebounds to 50% in 2018 then gradually declines to 28% by 2020. Topic 2 sees a modest resurgence from 26% in 2019 to 36% in 2020 (“Community & Artistic Practice” and “Threat” themes reemerge as the public grapples with automated creativity). In 2021, following the GPT-3.5 and Stable Diffusion releases, Topic 1 (“Utility & Adoption”) shoots up to 43%—its first significant appearance since 2013—while Topic 0 plummets below 10%. From 2022 onward, Topics 0 and 1 share roughly equal footing ($\sim$25–35% each), Topic 2/3 (“Community” and “Threat”) climb slowly but remain below 30%, and Topic 4 reappears variably (10–20%) around the EU AI Act negotiations and Andersen v. Stability AI lawsuit in 2023.
3.3 Readability Gradient
Figure 3: Readability across both axes. A: Readability across the Artistic to Technical Engineering discourse spectrum (D1). B: Readability across the General Market to Legal & Regulatory discourse spectrum (D2).
We computed Flesch–Kincaid grade levels for each 500-word section and plotted them against the consensus $D1$ (Artistic to Engineering) and $D2$ (Market to Legal) coordinates (Fig. 3). Sections in the top 5% of $D1$ and $D2$ (coloured crimson) consistently contain advanced word usage, usually requiring graduate and post-graduate level of education to understand. These sections are more than three grade-levels above the corpus median, whereas sections in the bottom 5% of each axis fall below or near the median readability. This confirms that our semantic axes are not just stylistic but also capture text complexity and register.
Across both axes we see a clear complexity gradient. Sections of text that lie at the high end (top 5%) of $D1$ and $D2$ tend to be written at a much higher Flesch–Kincaid grade-level than the bulk of the corpus. High-$D1$ passages (red dots on the right of plot A) are the furthest towards the technical engineering end of the spectrum. These passages almost all sit above the grade-level “cloud” of the rest of the data, suggesting that when discourse moves into detailed model architectures, training regimes, and other technical engineering descriptions it becomes substantially harder to read.
On the other hand, low-$D1$ passages (blue Xs on the left of plot A) tend to be below the median readability of the corpus, indicating those sections use more accessible, less jargon-filled language. Similar trends can be seen from high to low values on the $D2$ axis, which goes from legal and regulatory to general market and conversational discourse, respectively. Legal documents, lawsuit discussions, and policy analyses push the readability towards the “post-graduate” levels.
3.4 Statistical Alignment with Artist Survey
Table 3: Z‐tests comparing section‐level frame proportions ($\widehat{P}$) to Lovato et al. survey frame proportions ($P_{0}$).
Frame $\widehat{P}$ $P_{0}$ $z$ pp–value
Threat 0.256 0.271 –0.721 0.471
Utility 0.314 0.197 5.245 <0.001<0.001
Transparency 0.186 0.351 –8.798 <0.001<0.001
Ownership 0.244 0.181 3.036 0.002
3.5 Alignment with Artist Frames
Table 3 summarizes the average alignment of discourse sections with each of the four artist-identified frames: Threat, Utility, Transparency, and Ownership. It reports, for each frame, the corpus-wide mean alignment $\hat{P}$, the Lovato survey proportion $P_{0}$, the $z$-statistic, and two-tailed $p$-value. In every case except Threat, public discourse deviates significantly from artists’ own priorities (all $p\leq 0.002$). Specifically:
  • Threat does not have a significantly different representation ($\hat{P}=0.256$ vs. $P_{0}=0.271$, z=0.721z=-0.721, $p=0.471$).
  • Utility is over-emphasized compared to artists’ valuation ($\hat{P}=0.314$ vs. $P_{0}=0.197$, $z=5.245$, $p<0.001$).
  • Transparency is most severely under-emphasized ($\hat{P}=0.186$ vs. $P_{0}=0.351$, z=8.798z=-8.798, $p<0.001$).
  • Ownership is also over-emphasized ($\hat{P}=0.244$ vs. $P_{0}=0.181$, $z=3.036$, $p<0.002$).
3.6 Linguistic Metrics
Figure 4: Median lexical diversity by topic over time with key milestones in light gray.
Figure 4 shows each topic’s median section-level lexical diversity from 2013 through 2025, with key milestones indicated in light gray. A few things stand out:
  • Topic 3 (“Threat”) jumps sharply in 2016 ($\sim 0.60$) and again in 2019–2020 (peaking at $\sim 0.69$), suggesting that when conversations turn explicitly to AI as a threat, authors deploy a wider variety of expressions—perhaps to diagnose new risks and metaphors.
  • Topic 2 (“Community & Artistic Practice”) peaks in 2021 ($\sim 0.62$), coinciding with the post–Stable Diffusion surge in artist-facing blog posts and forum threads, which often mix technical how-to’s, personal testimony, and creative manifesto language.
  • In contrast, Topic 1 (“General Market”) and Topic 4 (“Legal & Regulatory”) gradually decline in diversity after 2017, bottoming out around 2023. That may reflect the standardization of market and policy jargon once key precedents and regulatory frameworks had been established.
Taken together, these oscillations in diversity map onto the technology and legal milestones we’ve already traced in Figure 2. Peaks in diversity often align with new narrative needs—e.g., fresh metaphors for threat (2016) and community responses (2021)—whereas more “settled” storylines around market sizing or legal norms show less lexical variety.
Figure 5: Correlations of linguistic and topic metrics (definitions in Table 3).
Metric Definition
Sentiment Transformer-based model (cardiffnlp/twitter-roberta-sentiment)
TextBlob Polarity TextBlob’s rule-based polarity score (1-1 to +1+1)
TextBlob Subjectivity TextBlob’s rule-based subjectivity score (0 to $1$)
Flesch-Kincaid Grade Textstat’s FK grade level readability metric
Lexical Diversity Unique tokens $\div$ total tokens per section
Log Article Length $\log_{e}(1+$ total word count of article))
D1 Axis Consensus UMAP-1 coordinate (Artistic to Technical)
D2 Axis Consensus UMAP-2 coordinate (Market to Legal)
Cluster Confidence Normalized distance-to-centroid pseudo-probability from BERTopic
Table 4: Metric definitions for Fig 5. Section-Level Correlations of Linguistic & Topic Metrics.
Figure 5 presents a Pearson-correlation heatmap over our section-level features (sentiment, subjectivity, readability, lexical diversity, log-article-word-count, $D1_{consensus}$, $D2_{consensus}$, and prob_consensus). A few key patterns emerge:
  • Lexical Diversity $\leftrightarrow$ Flesch–Kincaid Grade (r=0.61r=-0.61): As FK grade goes up, diversity goes down. In practice, the highest-grade passages (post-graduate readability) live in highly specialized subfields—either deep engineering or formal legal discourse—which tend to recycle a smaller core of technical or legal terms rather than drawing on a broad vocabulary.
Axes vs. Diversity & Readability
  • $D1_{consensus}\leftrightarrow$ Lexical Diversity (r=0.26r=-0.26): Moving toward the technical end of $D1$ correlates with lower diversity, echoing our readability finding that high-$D1$ sections demand higher grade-levels to parse (they use a narrow band of jargon repeatedly).
  • $D2_{consensus}\leftrightarrow$ Lexical Diversity ($r=+0.40$): Moving up the legal/IP end of $D2$ correlates with higher diversity—legal and policy debates often invoke a broader palette of statutory, philosophical, and economic terms.
Axes vs. Sentiment & Subjectivity
  • $D1_{consensus}\leftrightarrow$ Sentiment (r=0.15r=-0.15): Technical passages skew slightly more negative—perhaps reflecting critical or cautionary assessments of model limitations.
  • $D2_{consensus}\leftrightarrow$ Sentiment ($r=+0.22$) and (r=0.36r=-0.36) with polarity: Legal/regulatory sections trend less positive and less subjective overall, consistent with the formal, prescriptive tone of policy and copyright discourse.
Probability of Consensus (prob_consensus)
The lack of strong correlations between prob_consensus and any other metric ($|r|<0.10$) indicates that our consensus-based topic assignments are not simply a byproduct of section length, readability, or sentiment—they represent a distinct semantic signal.
4 Discussion
Our analysis uncovers a persistent gap between the dominant frames of public discourse on generative AI art and the lived priorities of artists themselves. By integrating large-scale, chronologically mapped media data with the survey benchmarks from Lovato et al. [9], we document both moments of alignment and misalignment that shed light on the cultural politics of AI in art.
4.1 Misalignment Between Public Discourse and Artists’ Frames
The corpus reveals that media and public discourse overwhelmingly foreground technical advancements (Topic 0: Technical engineering) and narratives of market utility and adoption (Topic 1), particularly following major AI breakthroughs. In contrast, the frames that artists themselves emphasize—especially transparency in model training data and concerns over ownership and labor rights—are consistently marginalized. The frames between the public and artists align mainly in response to episodic “flashpoint events”, such as landmark lawsuits (e.g., Andersen v. Stability AI) or major legislative debates (e.g., EU AI Act).
Direct hypothesis tests (Table 3) reinforce this divergence. Public discourse systematically underrepresents calls for transparency ($\hat{P}=0.186$ vs. $P_{0}=0.351$, $p<0.001$), while overemphasizing narratives of AI’s utility ($\hat{P}=0.314$ vs. $P_{0}=0.197$, $p<0.001$) and ownership debates ($\hat{P}=0.244$ vs. $P_{0}=0.181$, $p=0.002$). “Threat” as a frame achieves rough parity ($\hat{P}=0.256$ vs. $P_{0}=0.271$). However, the topic never dominates annual discourse, despite being the majority experience among surveyed artists. These patterns suggest that mainstream channels are structurally less responsive to artist-driven anxieties, except around moments of public controversy.
4.2 Dynamics and Structure of Discourse
Our multidimensional approach reveals that thematic dominance is neither stable nor symmetric. Technical engineering (Topic 0) frames surge and recede in tandem with new model releases, while legal and transparency concerns remain episodic. Legal and ethical (Topic 4) frames amplify around regulatory and artistic “flashpoints” but otherwise recede from view. Utility narratives, which have historically been amplified by market excitement and optimism, have grown increasingly prominent in the public conversation from 2020 onwards, vastly outpacing surveyed artists’ more measured or ambivalent stance on AI’s benefits.
Moreover, the consensus-based embedding and clustering pipeline surfaces not only five stable discourse frames, but also demonstrates that debates are structured along two continuous semantic axes—spanning from community-oriented, future-focused artistic to technical-engineering discourse ($D1$), and from general market to formal legal and regulatory discourse ($D2$). This multidimensionality captures the richness of public conversations, while clarifying where and how artists’ concerns are sidelined.
4.3 Linguistic Accessibility and Gatekeeping
A key, and often overlooked, equity finding is the pronounced gradient in linguistic complexity across discourse axes. Texts that score highly on the technical ($D1$) and legal ($D2$) spectra demand post-graduate reading levels, effectively raising barriers for artists and non-specialists. Legal and technical “gatekeeping” may not be intentional, but risks disempowering those whose economic livelihoods and creative identities are most affected by generative-AI. Instead, general-market and artistic frames trend towards greater readability, better aligning with practices of inclusive dialogue. Any effort to democratize policy or public engagement in this space must confront these structural hurdles to accessibility.
4.4 Diversity, Style, and the Evolution of Public Frames
Our longitudinal analysis of lexical diversity demonstrates that the texture of discourse is dynamic, peaking around narrative innovation and shrinking as market or legal vocabularies stabilize. For instance, the language of “Threat” reaches its expressive height when new risks are first debated. Further, legal and policy topics see declining diversity as regulatory schemas settle. This cycle may imply that public narratives calcify quickly, which could potentially shut down new ways of articulating artist concerns or imagining alternative futures as debates mature.
4.5 Theoretical Implications
By aligning machine-discovered topic clusters with survey-elicited artist priorities, our work advances both computational cultural analytics and the critical social study of AI in the arts. We show empirically how stakeholder frames may be dominated by technological and market logics, even as artists themselves remain divided, ambivalent, and resistant to simplistic narratives. The findings complicate the “pro-AI” vs. “anti-AI” binary, suggesting instead a rotating spotlight of public salience with artist-centered issues persistently struggling for sustained visibility.
4.6 Limitations
Our analysis is subject to several limitations. First, our search-term-based, English-only corpus is likely to under-sample non-Western, non-English voices, and may not capture intracultural variation in artistic priorities. Our methodology privileges prominent, indexed, or digitized content and may overlook vernacular or grassroots expressions. Finally, we are constrained by the coarseness of topic mapping. Future work would benefit from closer-genre-specific analysis and triangulation with further qualitative inquiry.
4.7 Future Directions
To deepen understanding and address these gaps, we propose three main expansions:
  1. 1.
    Sub-genre and Network Analysis: Disaggregate topic trends by media type and map influence networks to trace how artist concerns do or do not map to mainstream narratives.
  2. 2.
    Mixed-Methods Integration: Combine this consensus-based, transformer-driven mapping with in-depth interviews of artists to ground computational findings in lived experience.
  3. 3.
    Equity and Readability Interventions: Design and evaluate interventions aimed at translating dense technical discourse into accessible language to artists and advocates, in partnership with communities most affected by AI-generated art.
Finally, future work will further develop this reproducible methodology for longitudinal and multimodal corpora [12].
5 Conclusion
By integrating transformer embeddings, consensus clustering, and aligned multidimensional projections, we have mapped over a decade of generative AI discourse in art through five robust thematic clusters. Our analysis reveals a persistent and consequential divergence: while artists emphasize threats to creative labor, demands for transparency, concerns about ownership, and questions of utility, these frames are consistently overshadowed in public and scholarly narratives by technical innovation and market optimism. Legal, ethical, and labor-focused issues only become salient during episodic “flashpoints,” such as lawsuits or regulatory debates, before receding from view.
This misalignment matters. When the priorities of those most affected by generative AI—artists and creative workers—are sidelined, technological and regulatory trajectories are shaped by a selective and often self-reinforcing version of progress. Systems built on such narratives risk perpetuating inequity beneath the banner of innovation.
Our findings call for more sustained, inclusive, and critical engagement in both media and policy with the concerns of creative practitioners. Future research should broaden this consensus-based, multimodal approach to include non-English, non-Western, and underrepresented voices, and should combine large-scale computational mapping with qualitative, in-depth engagement with artists and communities. Only by bridging this discursive divide can the evolution of generative AI in art be guided by the diverse values, rights, and aspirations of those it most directly impacts.
Acknowledgments
We gratefully acknowledge the advice, support, and inspiration of many colleagues and collaborators throughout this project. We extend a special thank you to Juniper L. Lovato, Julia Zimmerman, and Matt Potts for their insight, encouragement, and generosity. Their contributions and guidance were invaluable to the development of this work.
Appendix A Appendix: Additional Tables & Figures
Search Phrase Relevant Documents in Top 10 Google Results
AI art 4
AI generated art 5
AI art impact on artists 7
Effects of AI image generators on artists’ careers 5
Artists’ responses to generative AI 7
Artists’ concerns about AI art 5
Interviews with artists about AI image generators 8
How AI affects freelance artists 9
AI copyright challenges for artists 12
Artists suing AI companies 10
How artists adapt to AI tools 6
Artists protest AI training data usage 7
Artists’ opinion on DeepDream art 10
AI art exhibition 8
Art incorporating AI 11
Creatives and AI 9
AI and the creative job market 11
Table 5: Breadth of search phrases and prevalence of relevant sources. Left: search phrases used to build the document corpus. Right: the number of relevant documents found in the first page of Google search results for each phrase. For some search terms, Google supplemented the standard 10 text links with up to 5 suggested video results, so the maximum possible total per search could exceed 10 (values above 10 reflect this).